OpenStudy (anonymous):

Gas Calculations Help!! Mass of magnesium Strip (grams): 0.032 g Volume of gas collected (mL): 30 mL Barometric pressure (atm): 1.1 atm Room temperature: 22C Vapor pressure of the water (torr): 19.8 torr

OpenStudy (anonymous):

1. Write the balanced equation for the reaction conducted in this lab, including appropriate phase symbols. Mg (s) + 2 H20 (l) --> H2 (g) + Mg (OH)2 (aq) 2. Determine the partial pressure of the hydrogen gas collected in the gas collection tube. 1.1atm = P1 + 0.0395atm 1.1atm – 0.0395atm = 1.0605 3. Calculate the moles of hydrogen gas collected. n = 1.0605atm * 30L / .0821 * 297 = 31.815 / 24.3837 = 1.1 mol 4. If magnesium was the limiting reactant in this lab, calculate the theoretical yield of the gaseous product. Show all steps of your calculation. 5. Determine the percent yield of this reaction, showing all steps of your calculation.

OpenStudy (anonymous):

Can you check my answers for 1-3 and help me with 4 and 5? Please and thank you

OpenStudy (jfraser):

1-3 looks good to me. If you know the Mg is the limiting reactant, then use stoichiometry to find the volume of H2 gas that SHOULD HAVE BEEN made. There's your theoretical yield

OpenStudy (anonymous):

Number 3 should be 1.3 mol instead of 1.1 mol

OpenStudy (jfraser):

did you convert your volume properly from mL to L?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@wolfe8 can you please help with this?

OpenStudy (australopithecus):

0.08235g/L is the density of H2 gas 30mL = 0.030L of H2 Gas (0.08235g/L) = x/0.030L where, x = grams of hydrogen gas once you have grams convert to moles using, Mole = grams/molecular mass

OpenStudy (australopithecus):

As for partial pressure I assume you are suppose to use, PV=nRT and solve for P so, P = nRT/V I got a partial pressure of about 1.08

OpenStudy (australopithecus):

4. If magnesium was the limiting reactant in this lab, calculate the theoretical yield of the gaseous product. Show all steps of your calculation. Use the moles of magnesium to find the moles of H2 gas, Mg (s) + 2 H20 (l) --> H2 (g) + Mg (OH)2 (aq) since it is 1 to 1 moles of hydrogen gas = moles of magnesium

OpenStudy (australopithecus):

Use the moles formula again I showed you above, that will give you the theoretical yeild

OpenStudy (australopithecus):

Percentage yield = (Actual Yield/Theoretical Yield)*100

OpenStudy (australopithecus):

Any questions?

OpenStudy (australopithecus):

For the gas constant R since you are dealing in atmospheres you will want to use the Gas constant with units, atm*L/K*mole so that you get atm for your units of pressure, saves you a conversion

OpenStudy (anonymous):

So for number 2 the partial pressure should be 1.08?

OpenStudy (australopithecus):

We are using the ideal gas model so it is fine.

OpenStudy (australopithecus):

You know volume of gas is 30mL, even if there was atmospheric pressure present it wouldn't matter. I'm crazy for thinking otherwise

OpenStudy (australopithecus):

Hence daltons law blah

OpenStudy (australopithecus):

Yeah, V = 0.030L T = 295K R = 0.0082 or something gas constant units with Latm n = Moles determined using volume

OpenStudy (australopithecus):

I'm making the assumption that the container size = the amount of hydrogen gas collected

OpenStudy (australopithecus):

https://en.wikipedia.org/wiki/Eudiometer

OpenStudy (anonymous):

I'm confused now.

OpenStudy (australopithecus):

Well if you can wait till later this afternoon I can try to dig up my old lab manual and help you with this, or you can search for calculations involving eudiometer, Try this, http://depts.gpc.edu/~dunchelb/1152L/Sampl_report.pdf Or type in google, eudiometer calculations

OpenStudy (australopithecus):

The atmospheric pressure effects the reading of the eudiometer so you have to use it in your calculation it is still ideal gas law but there is a step before it

OpenStudy (australopithecus):

This is a first year chem experiment though so I'm sure you can find out how to solve it properly, if there was no mention of an ediometer you can probably just negate the barometric pressure though and stick with the way I calculated it.

OpenStudy (australopithecus):

Hope this was helpful I need to get back to the grind

OpenStudy (australopithecus):

Oh and if you do it the eudiometer way you will not be converting volume to moles using density you will be using the ideal gas law to solve for moles

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

what was the answer?